Find the mean.
-9,-7,3,5,6,20

Found 2 solutions by apachemexican1, pcsp_kumat:
Answer by apachemexican1(64)

the mean of a set of numbers means you add all
the numbers together and divide the sum by the
amount of numbers you added together.
-9+(-7)=-16
-16+3=-13
-13+5=-8
-8+6=-2
-2+20=18
now divide the sum by 6
18/6=3
so the mean is 3.

Answer by pcsp_kumat(6)

mean=sum of values/no. of values
mean=[(-9)+(-7)+3+5+6+20]/6
=18/6
=3